# Distributed quantum computing with black-box subroutines

### v1

@@ I Introduction
The focus of our study is distributed quantum computing, which is believed to be a vital path to scale up quantum chips CAF24; BCD25. There are inspiring progresses in recent years, including distributed quantum algorithms ACR21; QLX24; TLW23; MGD23; Hwang_2025, techniques to break circuits into smaller blocks such as the circuit knitting PS23; UPR23; LMH23, and also experimental efforts to develop the interface among different types of qubits JYL25; AYH25. However, to the best of our knowledge, there are two important aspects which have not been well studied. First, compared with the usual circuit model describing a single chip, a massive amount of communication is required, and generic strategies to reduce the communication cost are less explored. Second, the remote control or execution of unknown subroutines is necessary and sometimes unavoidable, and this has not been studied for distributed quantum computing.
@@ V.2 Oblivious amplitude amplification
The OAA and also standard AA increase the circuit depth significantly by using iterative walk operators. This can be reduced to the original one (as $G$) by using OQT if we only need observable measurement on state $U|\psi\rangle$, instead of the state itself. The operators $G$, $GR$, and $G^{\dagger}R$ can be stored as Choi program states, and OQT on them can compose them together. This needs $O(\sqrt{p})$ program states, and this is quadratically better than an incoherent method, which measures the ancilla and repeat if a failure occurs.
@@ VI Physical realizations
In this section, we study the physical realizations of OQT, OQC, and others. We assess the main cost of obliviousness by comparing to their non-oblivious counterparts of the algorithms above, and layout the technical requirements on current quantum platforms.
@@ VI.2 Gates and qubits
Here we study how to realize our schemes above in currently available platforms, and we find there are notable differences. At present, there are a few leading quantum platforms with mainly two types of qubits:
@@
For photon-qubit platforms, a scalable approach is the cluster-state quantum computing RB01 rather than the circuit model due to the probabilistic nature of entangling gates KLM01. It realizes a gate sequence by a sequence of gate teleportation. The smallest cluster is just an ebit. For polarization or dual-rail encoded qubits, ebits are relatively easy to prepare, e.g., generated via SPDC scheme NC00 without using CNOT gates. Starting from ebits, a cluster state can be generated by the so-called fusion operations BR05. For hardware, the region for generation of cluster state and the region for implementation of measurements can be separated. These regions or chips can be connected by optical fibers Abu24, for instance. Photons can be transmitted directly or via teleportation, which is a merit of the photon-qubit platform.
@@ VII.3 Comparison with other protocols
First, let us recall our protocol in the bipartite case. The protocol, abbreviated as distributed black-box quantum computing (DBQC) in the Table, compute overlaps $|\langle\psi_{\text{o}}|U_{\text{o}}U_{\text{in}}|\psi_{\text{in}}\rangle|^{2}$ with a user Alice holding the input state $|\psi_{\text{in}}\rangle$, the classical description of it $[\psi_{\text{in}}]$, and the program state $|U_{\text{in}}\rangle$, and a server Bob holding the final measurement specified by a set of $\{|\psi_{\text{o}}\rangle\}$, their classical description $[\psi_{\text{o}}]$, and another program state $|U_{\text{o}}\rangle$.
This is directly distinct from the ‘obvious’ protocol when Alice knows $[U_{\text{in}}]$ and Bob knows $[U_{\text{o}}]$. This is the distributed quantum computing (DQC) in the Table. Actually, each participant can apply a gate $U_{i}$ directly and use QT to send its system to others. A nonlocal gate can be done remotely or belongs to a nonlocal participant. This setting is common and a few methods have been developed, such as distributed schemes for some quantum algorithms ACR21; QLX24; TLW23; MGD23; Hwang_2025 and the circuit knitting technique PS23; UPR23; LMH23. The obvious setting is easier due to the lack of obliviousness, and some techniques could also benefit DBQC. For instance, the circuit knitting aims to break down a nonlocal gate into a sum of local ones. An example is the expansion of a two-body gate in terms of Pauli gates
@@
OT $|\psi_{\text{in}}\rangle$, $[\psi_{\text{in}}]$, $|\psi_{\text{o}}\rangle$, $[\psi_{\text{o}}]$ $U$, $[U]$
BQC $|\psi_{\text{in}}\rangle$, $[\psi_{\text{in}}]$, $|\psi_{\text{o}}\rangle$, $[\psi_{\text{o}}]$, $[U]$ $U$
QvN $|\psi_{\text{in}}\rangle$, $[\psi_{\text{in}}]$, $|\psi_{\text{o}}\rangle$, $[\psi_{\text{o}}]$, $|U\rangle$ $U$, $[U]$
DBQC $|\psi_{\text{in}}\rangle$, $[\psi_{\text{in}}]$, $|U_{\text{in}}\rangle$ $|\psi_{\text{o}}\rangle$, $[\psi_{\text{o}}]$, $|U_{\text{o}}\rangle$
DQC $|\psi_{\text{in}}\rangle$, $[\psi_{\text{in}}]$, $|U_{\text{in}}\rangle$, $[U_{\text{in}}]$ $|\psi_{\text{o}}\rangle$, $[\psi_{\text{o}}]$, $|U_{\text{o}}\rangle$, $[U_{\text{o}}]$
@@ Appendix B Oblivious transfer and Bit commitment
It is well known that quantum capacity is private against eavesdroppers, while it often assumes both the sender Alice and the receiver Bob are honest. In a sense, the standard quantum communication is both committing and oblivious. Namely, for commitment, Alice commits and sends a state $|\psi\rangle$ directly. There is no way for Alice to change it anymore, neither for Bob to decode it until Alice tell him what it is, i.e., a classical description $[\psi]$ of it. For oblivious transfer, Alice and Bob first agreed upon an assignment $i\mapsto|\psi_{i}\rangle$, for a set of index $i$ that does not reveal $\psi_{i}$, and Bob only knows his index while Alice holds all the states without knowing them. Bob tells Alice the index $i$ and Alice then sends the state $|\psi_{i}\rangle$ to Bob. Although Alice knows $i$, she would not know the state $|\psi_{i}\rangle$. For both protocols, Bob can finally use $[\psi]$ and quantum verification schemes to verify the state $|\psi\rangle$. Our scheme for DBQC belongs to category (4), which does not allow entanglement assistance, hence the entanglement attack does not apply. However, it cannot tolerant dishonest participants who can use the wrong programs or send wrong measurement outcomes. Whether our protocol can be boosted towards secure multi-party quantum computation remains an interesting task.

### v2

@@ I Introduction
The focus of our study is distributed quantum computing, which is believed to be a vital path to scale up quantum chips CAF24; BCD25. There are inspiring progresses in recent years ACR21; QLX24; TLW23; MGD23; Hwang_2025; PS23; UPR23; LMH23; JYL25; AYH25; GPA24; AFM24; LZFD24, including distributed quantum algorithms ACR21; QLX24; TLW23; MGD23; Hwang_2025, techniques to break circuits into smaller blocks such as the circuit knitting PS23; UPR23; LMH23, and also experimental efforts to develop the interface among different types of qubits JYL25; AYH25. However, to the best of our knowledge, there are two important aspects which have not been well studied. First, compared with the usual circuit model describing a single chip, a massive amount of communication is required, and generic strategies to reduce the communication cost are less explored. Second, the remote control or execution of unknown subroutines is necessary and sometimes unavoidable, and this has not been studied for distributed quantum computing.
@@ V.2 Oblivious amplitude amplification
The OAA and also standard AA increase the circuit depth significantly by using iterative walk operators. This can be reduced to the original one (as $G$) by using OQT if we only need observable measurement on state $U|\psi\rangle$, instead of the state itself. The operators $G$, GRGR, and $G^{\dagger}R$ can be stored as Choi program states, and OQT on them can compose them together. This needs $O(\sqrt{p})$ program states, and this is quadratically better than an incoherent method, which measures the ancilla and repeat if a failure occurs.
@@ VI Physical realizations
In this section, we study the physical realizations of OQT, OQC, and others. All the tasks can be well described and understood as in the circuit model, such as the creation of program states and measurement on them. We assess the main cost of obliviousness by comparing to their non-oblivious counterparts of the algorithms above, and layout the technical requirements on current quantum platforms.
@@ VI.2 Gates and qubits
Here we study how to realize our schemes above in currently available platforms for universal quantum computing. We find that although being challenging, such as realizing Toffoli-type gates, these operations indeed can be realized, and there are notable differences across these platforms. At present, there are a few leading quantum platforms with mainly two types of qubits:
@@
For photon-qubit platforms, a scalable approach is the cluster-state quantum computing RB01 rather than the circuit model due to the probabilistic nature of entangling gates KLM01. It realizes a gate sequence by a sequence of gate teleportation. The smallest cluster is just an ebit. For polarization or dual-rail encoded qubits, ebits are relatively easy to prepare, e.g., generated via spontaneous parametric down-conversion (SPDC) scheme NC00 without using CNOT gates. Starting from ebits, a cluster state can be generated by the so-called fusion operations BR05. For hardware, the region for generation of cluster state and the region for implementation of measurements can be separated. These regions or chips can be connected by optical fibers Abu24, for instance. Photons can be transmitted directly or via teleportation, which is a merit of the photon-qubit platform.
@@ VII.3 Comparison with other protocols
First, let us recall our protocol in the bipartite case. The protocol, abbreviated as distributed black-box quantum computing (DBQC) in the Table, compute overlaps $|\langle\psi_{\text{o}}|U_{\text{o}}U_{\text{in}}|\psi_{\text{in}}\rangle|^{2}$ with a user Alice holding the input state $|\psi_{\text{in}}\rangle$, the classical description of it $\psi_{\text{in}}]$, and the program state $|U_{\text{in}}\rangle$, and a server Bob holding the final measurement specified by a set of $\{|\psi_{\text{o}}\rangle\}$, their classical description $\psi_{\text{o}}]$, and another program state $|U_{\text{o}}\rangle$.
This is directly distinct from the ‘obvious’ protocol when Alice knows $U_{\text{in}}]$ and Bob knows $U_{\text{o}}]$. This is the distributed quantum computing (DQC) in the Table. Actually, each participant can apply a gate $U_{i}$ directly and use QT to send its system to others. A nonlocal gate can be done remotely or belongs to a nonlocal participant. This setting is common and a few methods have been developed, such as distributed schemes for some quantum algorithms ACR21; QLX24; TLW23; MGD23; Hwang_2025 and the circuit knitting technique PS23; UPR23; LMH23. The obvious setting is easier due to the lack of obliviousness, and some techniques could also benefit DBQC. For instance, the circuit knitting aims to break down a nonlocal gate into a sum of local ones. An example is the expansion of a two-body gate in terms of Pauli gates
@@
OT $|\psi_{\text{in}}\rangle$, $\psi_{\text{in}}]$, $|\psi_{\text{o}}\rangle$, $\psi_{\text{o}}]$ $U$, $[U]$
BQC $|\psi_{\text{in}}\rangle$, $\psi_{\text{in}}]$, $|\psi_{\text{o}}\rangle$, $\psi_{\text{o}}]$, $[U]$ $U$
QvN $|\psi_{\text{in}}\rangle$, $\psi_{\text{in}}]$, $|\psi_{\text{o}}\rangle$, $\psi_{\text{o}}]$, $|U\rangle$ $U$, $[U]$
DBQC $|\psi_{\text{in}}\rangle$, $\psi_{\text{in}}]$, $|U_{\text{in}}\rangle$ $|\psi_{\text{o}}\rangle$, $\psi_{\text{o}}]$, $|U_{\text{o}}\rangle$
DQC $|\psi_{\text{in}}\rangle$, $\psi_{\text{in}}]$, $|U_{\text{in}}\rangle$, $U_{\text{in}}]$ $|\psi_{\text{o}}\rangle$, $\psi_{\text{o}}]$, $|U_{\text{o}}\rangle$, $U_{\text{o}}]$
@@ Appendix B Oblivious transfer and Bit commitment
It is well known that quantum capacity is private against eavesdroppers, while it often assumes both the sender Alice and the receiver Bob are honest. In a sense, the standard quantum communication is both committing and oblivious. Namely, for commitment, Alice commits and sends a state $|\psi\rangle$ directly. There is no way for Alice to change it anymore, neither for Bob to decode it until Alice tell him what it is, i.e., a classical description $\psi]$ of it. For oblivious transfer, Alice and Bob first agreed upon an assignment $i\mapsto|\psi_{i}\rangle$, for a set of index $i$ that does not reveal $\psi_{i}$, and Bob only knows his index while Alice holds all the states without knowing them. Bob tells Alice the index $i$ and Alice then sends the state $|\psi_{i}\rangle$ to Bob. Although Alice knows $i$, she would not know the state $|\psi_{i}\rangle$. For both protocols, Bob can finally use $\psi]$ and quantum verification schemes to verify the state $|\psi\rangle$. Our scheme for DBQC belongs to category (4), which does not allow entanglement assistance, hence the entanglement attack does not apply. However, it cannot tolerant dishonest participants who can use the wrong programs or send wrong measurement outcomes. Whether our protocol can be boosted towards secure multi-party quantum computation remains an interesting task.